# General Perturbative Expansion and Relations with Unitarity Triangles

### v1

@@ Abstract
In this letter, using a rephasing invariant formula $\delta=\arg[{V_{ud}V_{us}V_{cb}V_{tb}/V_{ub}\det V_{\rm CKM}}]$, we evaluate the CP phase $\delta$ of the CKM matrix $V_{\rm CKM}$ perturbatively for small quark mixing angles $s_{ij}^{u,d}$ with associated phases $\rho_{ij}^{u,d}$. Consequently, we derived a relation $\delta\simeq\arg[\Delta s_{12}\Delta s_{23}/(\Delta s_{13}-s^{u}_{12}e^{-i\rho^{u}_{12}}\Delta s_{23})]$ with $\Delta s_{ij}\equiv s^{d}_{ij}e^{-i\rho^{d}_{ij}}-s^{u}_{ij}e^{-i\rho^{u}_{ij}}$. Such a result represents the analytic behavior of the CKM phase. The uncertainty in the relation is of order $O(\lambda^{2})\sim 4\%$, which is comparable to the current experimental precision. Comparisons with experimental data suggest that the hypothesis of some CP phases being maximal. We also discussed relationships between the phase $\delta$ and unitarity triangles. As a result, several relations between the angles $\alpha,\beta,\gamma$ and $\delta$ are identified through other invariants $V_{il}V_{jm}V_{kn}/\det V_{\rm CKM}$.
@@ I Introduction
Understanding the origin of CP violation is crucial for explaining the baryon asymmetry in the universe. CP violation in the CKM matrix Kobayashi:1973fv has been discussed in various forms throughout the history of particle physics Fritzsch:1995nx ; Rasin:1997pn ; Harrison:2009bz ; Xing:2020ijf . In particular, Ref. Hall:1993ni presented a general treatment of the CKM matrix and its CP-violating phase. However, such general treatments have not been pursued in later literature. Furthermore, most analyses often use the Jarlskog invariant Jarlskog:1985ht , which is of order $O(10^{-5})$ in the quark sector. Since the small invariant is highly sensitive to various approximations, capturing its analytic behavior—including error estimation—has been a technically challenging task from the standpoint of computational precision.
In this letter, employing a recently proposed rephasing invariant formula Yang:2025hex , we derive a general perturbative expression for the CP-violating phase in the CKM matrix, and its phenomenological consequences.
II A Rephasing Invariant formula for CP Phase of CKM Matrix
and Its Perturbative Expansion
We begin by presenting a method to directly extract the CP phase $\delta$ from the CKM matrix $V_{\rm CKM}$ defined in a general phase basis. To convert a given $V_{\rm CKM}$ into the PDG standard parametrization $V_{\rm CKM}^{0}$, we remove unphysical phases by applying redefinition of phases as
@@ II A Rephasing Invariant formula for CP Phase of CKM Matrix
Here, $\Psi_{L,R}={\rm diag}(e^{i\gamma_{(L,R)1}},\,e^{i\gamma_{(L,R)2}},\,e^{i\gamma_{(L,R)3}})$ are diagonal phase matrices, and each $\gamma_{(L,R)i}$ represents an arbitrary phase. Due to the overall phase redundancy, the number of independent degrees of freedom is five.
In the standard PDG parametrization, elements of the mixing matrix satisfy the following conditions:
$\arg V_{ud}^{0}=\arg V_{us}^{0}=\arg V_{cb}^{0}=\arg V_{tb}^{0}=0\,,~~\arg\left[V_{cd}^{0}V_{ts}^{0}-V_{cs}^{0}V_{td}^{0}\right]=\arg\left[V_{ub}^{0*}\det V_{\rm CKM}^{0}\right]=\delta\,.$ (2)
The final constraint follows from the identity for the inverse of a unitary matrix.
Using these five conditions, one can solve for the five phase parameters $\gamma_{(L,R)i}$. As a result, for a CKM matrix given in an arbitrary basis of phases, the CP-violating phase $\delta$ in the PDG convention is expressed as
$\delta$ $=\arg\left[\frac{V_{ud}V_{us}V_{cb}V_{tb}}{V_{ub}\det V_{\rm CKM}}\right]\,.$ (3)
This expression is explicitly rephasing invariant, including the phase of $\det V_{\rm CKM}$ Wu:1985ea ; Bjorken:1987tr ; Nieves:1987pp ; Jenkins:2007ip , and it clearly coincides with the phase $\delta$ in the PDG standard parametrization.
@@ 1. Factorizability:
It can be decomposed into individual elements $V_{\alpha\beta}$ and the determinant $\det V_{\rm CKM}$, making the computation more transparent and straightforward.
2. Robustness under approximations:
The Jarlskog invariant $J$ is highly sensitive to approximations because of its smallness $O(10^{-5})$ in the quark sector. On the other hand, the new invariant is of order $O(1)$ and therefore less affected by perturbative corrections. Furthermore, its uncertainty is easier to quantify.
@@ 3. Completeness of phase information:
We now demonstrate the relation between this formula and the well-known Jarlskog invariant. By dividing the complex quantity (inside the argument) by its modulus, the phase is explicitly extracted as
@@
The mixing angles $s_{ij}^{u,d}\equiv\sin\theta_{ij}^{u,d}\,,\,c_{ij}^{u,d}\equiv\cos\theta_{ij}^{u,d}$ of $U_{u,d}$ are assumed to satisfy $s_{12}^{u,d}\sim\lambda$, $s_{23}^{u,d}\sim\lambda^{2}$, and $s_{13}^{u,d}\sim\lambda^{3}$, with an expansion parameter $\lambda\simeq 0.2$.
@@
We now proceed to define the notation of perturbative expansion. The matrices $U_{u,d}$ are generally written as $U_{u,d}=\Phi_{u,d}^{L}U_{u,d}^{0}\Phi_{u,d}^{R}$ with diagonal phase matrices $\Phi_{u,d}^{L,R}$ and their PDG parametrizations $U_{u,d}^{0}$. Due to the freedom of right-handed phase transformations, the unitary matrices are redefined as $U_{u,d}^{1}=\Phi_{u,d}^{L}U_{u,d}^{0}\Phi_{u,d}^{L\dagger}$ without loss of generality. Since $U_{u,d}^{0}$ contain only small mixing angles, the leading order of the perturbation is approximated as follows:
$U_{u,d}^{1}\simeq\begin{pmatrix}1&s_{12}^{u,d}e^{-i\rho_{12}^{u,d}}&s_{13}^{u,d}e^{-i\rho_{13}^{u,d}}\\ -s_{12}^{u,d}e^{i\rho_{12}^{u,d}}&1&s_{23}^{u,d}e^{-i\rho_{23}^{u,d}}\\ -s_{13}^{u,d}e^{i\rho_{13}^{u,d}}+s_{12}^{u,d}s_{23}^{u,d}e^{i\rho_{12}^{u,d}+i\rho_{23}^{u,d}}&-s_{23}^{u,d}e^{i\rho_{23}^{u,d}}&1\\ \end{pmatrix},$ (7)
where $\rho_{ij}^{u,d}$ are the associated CP-violating phases corresponding to the mixing angles. The next-to-leading order terms in each matrix element are suppressed by at least order $\lambda^{2}$ compared to the leading order. Since the right-handed phases $\Phi^{R}_{u,d}$ of quarks do not affect the observed CP phase, we will omit them hereafter.
@@
Since errors of this expression are $O(\lambda^{2})\simeq 4\%$, which is comparable to the current experimental uncertainties, the expression is sufficiently accurate. In particular, the absolute value of the denominator has a fixed value,
@@
$\left|V_{us}\over V_{ub}/V_{cb}\right|e^{i\delta}$ $\simeq\frac{s^{d}_{12}e^{-i\rho^{d}_{12}}-s^{u}_{12}e^{-i\rho^{u}_{12}}}{-s^{u}_{12}e^{-i\rho^{u}_{12}}}+O(\lambda^{2})\,,$ (15)
@@
The observed physical value of CKM matrix in the latest UTfit is UTfit:2022hsi
$\sin\theta_{12}^{\rm CKM}$ $=0.22519\pm 0.00083\,,~~~\sin\theta_{23}^{\rm CKM}=0.04200\pm 0.00047\,,$ (17)
$\sin\theta_{13}^{\rm CKM}$ $=0.003714\pm 0.000092\,,~~~\delta_{q}=1.137\pm 0.022\,.$ (18)
Given this experimental value, the hypothesis that the phase is maximal $\rho^{d}_{12}-\rho^{u}_{12}=\pi/2$ appears quite plausible Shin:1985cg ; Gronau:1985tx ; Fritzsch:1985yv ; Kang:1985nw ; Lehmann:1995br ; Kang:1997uv ; Fritzsch:1999im ; Antusch:2009hq ; Xing:2003yj ; Yang:2020qsa ; Yang:2020goc ; Yang:2021smh . A non-perturbative treatments beyond perturbation theory for $s_{12}^{f}$ and $s_{23}^{f}$ are also available in previous papers Yang:2024ulq ; Yang:2025yst .
@@
It is of particular interest to discuss relations between the results and unitarity triangles Harrison:2006bj ; Frampton:2010ii ; Dueck:2010fa ; Li:2010ae . From the alternative $s$-$b$ unitarity triangle, given by $V_{us}V_{ub}^{*}+V_{cs}V_{cb}^{*}+V_{ts}V_{tb}^{*}=0$, its three angles are rephasing invariant quantities,
$\alpha^{\prime}=\arg\left[-{V_{ts}V_{tb}^{*}\over V_{us}V_{ub}^{*}}\right],~~\beta^{\prime}=\arg\left[-{V_{cs}V_{cb}^{*}\over V_{ts}V_{tb}^{*}}\right],~~\gamma^{\prime}=\arg\left[-{V_{us}V_{ub}^{*}\over V_{cs}V_{cb}^{*}}\right].$ (19)
@@
$\gamma^{\prime}+\pi=\arg\left[{V_{us}V_{cb}\over V_{ub}V_{cs}}\right]\simeq\delta\,.$ (20)
@@
$\delta+\alpha^{\prime}=\arg\left[\frac{V_{ud}V_{us}V_{cb}V_{tb}}{V_{ub}\det V_{\rm CKM}}\right]+\arg\left[-{V_{ts}V_{tb}^{*}\over V_{us}V_{ub}^{*}}\right]=\arg\left[-\frac{V_{ud}V_{cb}V_{ts}}{\det V_{\rm CKM}}\right]=1.05^{\circ}\,,$ (21)
$\delta-\gamma^{\prime}-\pi=\arg\left[\frac{V_{ud}V_{us}V_{cb}V_{tb}}{V_{ub}\det V_{\rm CKM}}\right]-\arg\left[{V_{cb}V_{us}\over V_{ub}V_{cs}}\right]=\arg\left[\frac{V_{ud}V_{cs}V_{tb}}{\det V_{\rm CKM}}\right]=-0.0019^{\circ}\,,$ (22)
@@
$\delta+\alpha+\pi$ $=\arg\left[\frac{V_{ud}V_{us}V_{cb}V_{tb}}{V_{ub}\det V_{\rm CKM}}\right]+\arg\left[{V_{td}V_{tb}^{*}\over V_{ud}V_{ub}^{*}}\right]=\arg\left[\frac{V_{us}V_{cb}V_{td}}{\det V_{\rm CKM}}\right]=-22.45^{\circ}\fallingdotseq-\beta\,,$ (24)
@@
Since the perturbative expansions of $V_{us}$ and $V_{cd}$ agree up to $\mathcal{O}(\lambda^{3})$ in Eq. (9), the phase of $-V_{us}V_{cd}V_{tb}$ is of order $\mathcal{O}(\lambda^{4})\sim 0.2\%$. At this point, it seems meaningful to examine the remaining two invariants;
$\arg\left[\frac{V_{ub}V_{cs}V_{td}}{\det V_{\rm CKM}}\right]=-87.60^{\circ}\fallingdotseq\alpha-\pi\,,~~~\arg\left[\frac{V_{us}V_{cd}V_{tb}}{\det V_{\rm CKM}}\right]=-64.01^{\circ}\simeq-\gamma\,.$ (27)
That is, three of the six invariants are approximately identified with the angles $\alpha-\pi$, $-\beta$, and $-\gamma$, and the remaining three represent small differences between the angles and invariants. Therefore, they provide an alternative perspective on the characteristic CP phases in the CKM matrix. A similar argument can be applied to other unitarity triangles.

### v2

@@ Abstract
In this letter, using a rephasing invariant formula $\delta=\arg[{V_{ud}V_{us}V_{cb}V_{tb}/V_{ub}\det V_{\rm CKM}}]$, we evaluate the CP phase $\delta$ in the CKM matrix $V_{\rm CKM}$ perturbatively for small quark mixing angles $s_{ij}^{u,d}$ with associated phases $\rho_{ij}^{u,d}$. Consequently, we derived a relation $\delta\simeq\arg[\Delta s_{12}\Delta s_{23}/(\Delta s_{13}-s^{u}_{12}e^{-i\rho^{u}_{12}}\Delta s_{23})]$ with $\Delta s_{ij}\equiv s^{d}_{ij}e^{-i\rho^{d}_{ij}}-s^{u}_{ij}e^{-i\rho^{u}_{ij}}$, which represents the analytic behavior of the CKM phase. The uncertainty in the relation is of order $O(\lambda^{2})\sim 4\%$, comparable to the current experimental precision. Experimental data suggest that the hypothesis of some CP phases being maximal. We also discussed relationships between the phase $\delta$ and unitarity triangles. As a result, several relations between the angles $\alpha,\beta,\gamma$ and $\delta$ are identified through other invariants $V_{il}V_{jm}V_{kn}/\det V_{\rm CKM}$.
@@ I Introduction
Understanding the origin of CP violation is crucial for explaining the baryon asymmetry in the universe. In the Standard Model, the CP violation in the quark sector is described by a single phase contained in the Cabibbo–Kobayashi–Maskawa (CKM) matrix Cabibbo:1963yz ; Kobayashi:1973fv . Although the mixing matrix depends on the choice of parameterization, physical observables must be formulated in terms of rephasing invariants Wu:1985ea ; Bjorken:1987tr ; Nieves:1987pp ; Branco:1999fs ; Kuo:2005pf ; Jenkins:2007ip ; Chiu:2012uc ; Chiu:2015ega .
Traditionally, the Jarlskog invariant $J$ Jarlskog:1985ht has played a central role in the description of CP violation. In the discussion of weak basis invariants (WBI) Bernabeu:1986fc ; Gronau:1986xb ; Botella:1994cs , it has been shown that the lowest-dimensional CP-odd WBI is proportional to $J$. However, since the invariant $J$ is a small quantity of order $O(10^{-5})$, it is highly sensitive to various approximations. Capturing analytic behavior of the CP violation—including error estimation—has been a technically challenging task due to the computational precision. Although Ref. Hall:1993ni presented a general perturbative treatment of the CKM matrix and its CP phase, such a general treatment has not been pursued in later literature.
In recent years, a rephasing invariant involving the determinant of the mixing matrix has been proposed for a direct evaluation of the CP phase Yang:2025hex . It complements conventional approaches in both of analytic structure and error estimation. In this letter, employing the rephasing invariant formula $\delta=\arg[{V_{ud}V_{us}V_{cb}V_{tb}/V_{ub}\det V_{\rm CKM}}]$, we derive a general perturbative expression for the CP phase in the CKM matrix, and its phenomenological consequences. This allows for precise analyses that were difficult to achieve with the Jarlskog invariant, which is a cubicly suppressed quantity by the three mixing angles. Furthermore, the insights from the quark sector are applicable to the analysis of the leptonic CP phase $\delta_{l}$. With forthcoming experiments, DUNE DUNE:2020jqi and T2HK Hyper-KamiokandeProto-:2015xww , such analyses are anticipated to shed light on underlying grand unified relations.
II A Rephasing Invariant formula for CP Phase of CKM Matrix and Its Perturbative Expansion
We begin by presenting a method to directly extract the CP phase $\delta$ from the CKM matrix $V_{\rm CKM}$ defined in a general phase basis. To convert a given $V_{\rm CKM}$ into the PDG standard parametrization $V_{\rm CKM}^{0}$, unphysical phases are removed by redefinition of phases as
@@ II A Rephasing Invariant formula for CP Phase of CKM Matrix and Its Perturbative Expansion
Here, $\Psi_{L,R}={\rm diag}(e^{i\gamma_{(L,R)1}},\,e^{i\gamma_{(L,R)2}},\,e^{i\gamma_{(L,R)3}})$ are diagonal phase matrices, and each $\gamma_{(L,R)i}$ represents phases to be solved. The inverse transformation from the PDG parametrization to the original basis $V_{\rm CKM}=\Psi_{L}V_{\rm CKM}^{0}\Psi_{R}^{\dagger}$ is written by
$\begin{pmatrix}V_{ud}&V_{us}&V_{ub}\\ V_{cd}&V_{cs}&V_{cb}\\ V_{td}&V_{ts}&V_{tb}\\ \end{pmatrix}=\begin{pmatrix}e^{i\gamma_{L1}}&0&0\\ 0&e^{i\gamma_{L2}}&0\\ 0&0&e^{i\gamma_{L3}}\\ \end{pmatrix}\begin{pmatrix}|V_{ud}|&|V_{us}|&|V_{ub}|e^{-i\delta}\\ V_{cd}^{0}&V_{cs}^{0}&|V_{cb}|\\ V_{td}^{0}&V_{ts}^{0}&|V_{tb}|\\ \end{pmatrix}\begin{pmatrix}e^{-i\gamma_{R1}}&0&0\\ 0&e^{-i\gamma_{R2}}&0\\ 0&0&e^{-i\gamma_{R3}}\\ \end{pmatrix}$. (2)
In particular, focusing on $V_{ub}=|V_{ub}|e^{-i(\delta-\gamma_{L1}+\gamma_{R3})}$,
$V_{ub}^{*}=|V_{ub}|e^{i(\delta-\gamma_{L1}+\gamma_{R3})}\,,~~V_{ub}^{*}e^{i(\gamma_{L1}-\gamma_{R3})}=|V_{ub}|e^{i\delta}\,.$ (3)
By using $\det V_{\rm CKM}=e^{i(\gamma_{L1}+\gamma_{L2}+\gamma_{L3}-\gamma_{R1}-\gamma_{R2}-\gamma_{R3})}$, the remaining four phases are eliminated as
$\arg\left[{V_{ud}V_{us}V_{cb}V_{tb}\over\det V_{\rm CKM}}\right]=\gamma_{L1}-\gamma_{R3}\,,~~\Rightarrow~~\delta=\arg\left[{V_{ud}V_{us}V_{cb}V_{tb}\over V_{ub}\det V_{\rm CKM}}\right]\,.$ (4)
Including the phase of $\det V_{\rm CKM}$, this formula is explicitly rephasing invariant, and it clearly coincides with the phase $\delta$ in the PDG parametrization. Except for this particular combination, the phase $\delta$ cannot be solved explicitly because it remains on the right-hand side. The subsequent papers Yang:2025ftl ; Yang:2025vrs show that other similar combinations correspond to CP phases in other parametrizations of the mixing matrix.
@@ 1. Factorizability:
The CP phase is decomposed into individual elements $V_{\alpha i}$ and the determinant $\det V_{\rm CKM}$, making the computation more transparent and straightforward. Furthermore, we can easily quantify uncertainties and errors of the phase. Although the Jarlskog invariant $J$ is highly sensitive to approximations because of its smallness, the new invariant is of order $O(1)$ and therefore less affected by perturbative corrections.
2. Independence of constraints:
The invariant $J$ simultaneously constrains the mixing angles and the phases. In this formula, since the phase depends only on arguments of matrix elements, the constraint is independent of those of mixing angles.
@@ 3. Completeness of phase information:
We demonstrate the relation between this formula and the well-known Jarlskog invariant. By dividing the complex quantity (inside the argument) by its modulus, the phase is reconstructed as
@@
The mixing angles $s_{ij}^{u,d}\equiv\sin\theta_{ij}^{u,d}\,,\,c_{ij}^{u,d}\equiv\cos\theta_{ij}^{u,d}$ of $U_{u,d}$ satisfy $s_{12}^{u,d}\sim\lambda$, $s_{23}^{u,d}\sim\lambda^{2}$, and $s_{13}^{u,d}\sim\lambda^{3}$, with an expansion parameter $\lambda\simeq 0.2$.
@@
We now proceed to define the notation of perturbative expansion. The matrices $U_{u,d}$ are generally written as $U_{u,d}=\Phi_{u,d}^{L}U_{u,d}^{0}\Phi_{u,d}^{R}$ with diagonal phase matrices $\Phi_{u,d}^{L,R}$ and their PDG parametrizations $U_{u,d}^{0}$. Due to the freedom of right-handed phase transformations, the unitary matrices are redefined as $U_{u,d}^{1}=\Phi_{u,d}^{L}U_{u,d}^{0}\Phi_{u,d}^{L\dagger}$. Since $U_{u,d}^{1}$ has a unit determinant, it is written as
$U_{u,d}^{1}=\begin{pmatrix}1&0&0\\ 0&c_{23}^{u,d}&s_{23}^{u,d}e^{-i\rho_{23}^{u,d}}\\ 0&-s_{23}^{u,d}e^{i\rho_{23}^{u,d}}&c_{23}^{u,d}\\ \end{pmatrix}\begin{pmatrix}c_{13}^{u,d}&0&s_{13}^{u,d}e^{-i\rho_{13}^{u,d}}\\ 0&1&0\\ -s_{13}^{u,d}e^{i\rho_{13}^{u,d}}&0&c_{13}^{u,d}\\ \end{pmatrix}\begin{pmatrix}c_{12}^{u,d}&s_{12}^{u,d}e^{-i\rho_{12}^{u,d}}&0\\ -s_{12}^{u,d}e^{i\rho_{12}^{u,d}}&c_{12}^{u,d}&0\\ 0&0&1\\ \end{pmatrix}$, (8)
where $\rho_{ij}^{u,d}$ are the associated CP-violating phases corresponding to the mixing angles. Under this setup, the leading order of the mixing matrix is approximated as
$U_{u,d}^{1}\simeq\begin{pmatrix}1&s_{12}^{u,d}e^{-i\rho_{12}^{u,d}}&s_{13}^{u,d}e^{-i\rho_{13}^{u,d}}\\ -s_{12}^{u,d}e^{i\rho_{12}^{u,d}}&1&s_{23}^{u,d}e^{-i\rho_{23}^{u,d}}\\ -s_{13}^{u,d}e^{i\rho_{13}^{u,d}}+s_{12}^{u,d}s_{23}^{u,d}e^{i\rho_{12}^{u,d}+i\rho_{23}^{u,d}}&-s_{23}^{u,d}e^{i\rho_{23}^{u,d}}&1\\ \end{pmatrix}$. (9)
The next-to-leading order terms in each matrix element are suppressed by at least order $\lambda^{2}$ compared to the leading order. Since the right-handed phases $\Phi^{R}_{u,d}$ of quarks do not affect the observed CP phase, we will omit them hereafter.
@@
The errors of this expression are $O(\lambda^{2})\simeq 4\%$, which is comparable to the current experimental uncertainties. In particular, the absolute value of the denominator has a fixed value,
@@
$\left|V_{us}\over V_{ub}/V_{cb}\right|e^{i\delta}$ $\simeq\frac{s^{d}_{12}e^{-i\rho^{d}_{12}}-s^{u}_{12}e^{-i\rho^{u}_{12}}}{-s^{u}_{12}e^{-i\rho^{u}_{12}}}+O(\lambda^{2})\,$,
@@
The observed physical value of the CKM matrix in the latest UTfit is UTfit:2022hsi
$\sin\theta_{12}^{\rm CKM}$ $=0.22519\pm 0.00083\,,~~~\sin\theta_{23}^{\rm CKM}=0.04200\pm 0.00047\,$,
$\sin\theta_{13}^{\rm CKM}$ $=0.003714\pm 0.000092\,,~~~\delta=1.137\pm 0.022=65.15\pm 1.3^{\circ}\,.$ (17)
Given this experimental value, the hypothesis that the phase is maximal $\rho^{d}_{12}-\rho^{u}_{12}=\pi/2$ appears quite plausible Shin:1985cg ; Gronau:1985tx ; Fritzsch:1985yv ; Kang:1985nw ; Lehmann:1995br ; Kang:1997uv ; Fritzsch:1999im ; Antusch:2009hq ; Xing:2003yj ; Yang:2020qsa ; Yang:2020goc ; Yang:2021smh .
By combining these results with the previous work, predictions of grand unified relations become much more straightforward. The general perturbative evaluation of the leptonic CP phase $\delta_{l}$ is approximately Yang:2025hex
$\delta_{l}\simeq\delta_{\nu}+2s^{e}_{23}\sin\rho_{23}^{e}+4.66s^{e}_{12}\sin(\delta_{\nu}-\rho_{12}^{e})+4.66s^{e}_{13}\sin(\delta_{\nu}-\rho_{13}^{e})\,$, (18)
where $\delta_{\nu}$ is the CP phase in the PDG parametrization of the neutrino diagonalization matrix $U_{\nu}$. The mixing angles $s^{e}_{ij}$ and CP phases $\rho^{e}_{ij}$ for the charged leptons are defined in the same way as for the quarks. If a large 1-2 mixing $s^{e}_{12}\sim 0.1$ and the associated phase $\rho_{12}^{e}\sim\pi/2$ exist as in the quark sector, they can contribute approximately $0.5\sim 30^{\circ}$ to $\delta_{l}$. Therefore, without accidental cancellations, the CP phase $\delta_{l}$ is expected to be discovered within roughly ten years of operations of DUNE DUNE:2020jqi and T2HK Hyper-KamiokandeProto-:2015xww . Non-perturbative treatments beyond perturbation theory for $s_{12}^{f}$ and $s_{23}^{f}$ are also available in previous papers Yang:2024ulq ; Yang:2025yst .
@@
It is of particular interest to discuss relations between the results and unitarity triangles Hocker:2006xb ; Frampton:2010ii ; Dueck:2010fa ; Li:2010ae . From the alternative $b$-$s$ unitarity triangle, given by $V_{ub}V_{us}^{*}+V_{cb}V_{cs}^{*}+V_{tb}V_{ts}^{*}=0$, its three angles are rephasing invariant quantities Wu:1994di ; Harrison:2006bj ; Harrison:2009bz ,
$\alpha^{\prime}=\arg\left[-{V_{tb}V_{ts}^{*}\over V_{ub}V_{us}^{*}}\right]=64.09^{\circ},~~\beta^{\prime}=\arg\left[-{V_{cb}V_{cs}^{*}\over V_{tb}V_{ts}^{*}}\right]=1.05^{\circ},~~\gamma^{\prime}=\arg\left[-{V_{ub}V_{us}^{*}\over V_{cb}V_{cs}^{*}}\right]=114.86^{\circ}.$ (19)
@@
$\pi-\gamma^{\prime}=\arg\left[{V_{us}V_{cb}\over V_{ub}V_{cs}}\right]=65.15^{\circ}\simeq\delta\,.$ (20)
@@
$\delta-\alpha^{\prime}=\arg\left[\frac{V_{ud}V_{us}V_{cb}V_{tb}}{V_{ub}\det V_{\rm CKM}}\right]-\arg\left[-{V_{tb}V_{ts}^{*}\over V_{ub}V_{us}^{*}}\right]=\arg\left[-\frac{V_{ud}V_{cb}V_{ts}}{\det V_{\rm CKM}}\right]=1.05^{\circ}\,$,
$\delta+\gamma^{\prime}-\pi=\arg\left[\frac{V_{ud}V_{us}V_{cb}V_{tb}}{V_{ub}\det V_{\rm CKM}}\right]+\arg\left[{V_{ub}V_{cs}\over V_{cb}V_{us}}\right]=\arg\left[\frac{V_{ud}V_{cs}V_{tb}}{\det V_{\rm CKM}}\right]=-0.0019^{\circ}\,$, (21)
@@
$\delta+\alpha+\pi$ $=\arg\left[\frac{V_{ud}V_{us}V_{cb}V_{tb}}{V_{ub}\det V_{\rm CKM}}\right]+\arg\left[{V_{td}V_{tb}^{*}\over V_{ud}V_{ub}^{*}}\right]=\arg\left[\frac{V_{us}V_{cb}V_{td}}{\det V_{\rm CKM}}\right]=-22.45^{\circ}\fallingdotseq-\beta\,$,
@@
Since the perturbative expansions of $V_{us}$ and $V_{cd}$ agree up to $\mathcal{O}(\lambda^{3})$ in Eq. (11), the phase of $V_{us}V_{cd}V_{tb}$ is of order $\mathcal{O}(\lambda^{4})\sim 0.2\%$. At this point, it seems meaningful to examine the remaining two invariants;
$\arg\left[-\frac{V_{ub}V_{cs}V_{td}}{\det V_{\rm CKM}}\right]=92.40^{\circ}\fallingdotseq\alpha\,,~~~\arg\left[\frac{V_{ub}V_{cd}V_{ts}}{\det V_{\rm CKM}}\right]=-64.06^{\circ}\simeq-\gamma\,.$ (25)
That is, three of the six invariants are approximately identified with the angles $\alpha-\pi$, $\beta$, and $\gamma$, and the remaining three represent small differences between the angles and invariants. Although such third-order invariants have already been discussed in the form fixed by $\det V=1$ Kuo:2005pf , this scheme provides an alternative perspective on the characteristic CP phases in the CKM matrix.